OpenStudy (anonymous):

how to find log of log50/15?

OpenStudy (anonymous):

@lgbasallote @Hero @Libniz @TranceNova

hartnn (hartnn):

is it (log 50)/15 or log(50/15) ?

OpenStudy (anonymous):

\[\log_{50/15} \]

hartnn (hartnn):

that has no meaning the base is 50/15 ?

OpenStudy (anonymous):

convert 50=5*10 and 15=5*3 so now =log10/3 =log10-log3 =1-log3

OpenStudy (anonymous):

now it easily solving i think

hartnn (hartnn):

lol, no...and she also needs to find log of that.....

OpenStudy (anonymous):

waht shold be the value of = -2.303*2*298*log50/15 ? now please elaboarate.

OpenStudy (anonymous):

i think -717.69

hartnn (hartnn):

can't u use calculator ? u can find log (50/15) = log(10/3) using calcy

OpenStudy (anonymous):

@ruchi k now can i ask 1 que.?

hartnn (hartnn):

* log(10/3) using calcy ONLY, not manually

OpenStudy (anonymous):

yah @vipul92 hey @hartnn btw the value is -1436 calories.

OpenStudy (anonymous):

@hartnn it's ok @vipul92 is just posting answers to people's posts and usually posting wrong answers

hartnn (hartnn):

-1436 is double the value of expression that u wrote

OpenStudy (anonymous):

hey @gaara you don't know how to multiply and subtract two no. and said that i m wrong

OpenStudy (anonymous):

excuse me.! @vipul92 this place is for posting questings and answering them, so please don't do useless talks here. that would go against the code of conduct! thanks!

OpenStudy (anonymous):

k as u wish

hartnn (hartnn):

using calculator -2.303*2*298*log(50/15) = -717.697 -1436 is double of that

OpenStudy (anonymous):

sorry to everone the question is something like this : -2.303*2*2*298*log(50/15) =

hartnn (hartnn):

-2.303*2*2*298*log(50/15)=-1436 using calculator

OpenStudy (anonymous):

hey i'm asking that how u hav find log in this question

hartnn (hartnn):

doesn't your calculator have the option to calculate log? if not then u can simplify log(50/15) = log (10/3) = log 10 -log 3 =1-log 3 and u need to know log 3 = 0.477

OpenStudy (anonymous):

in xamination calculator r nt preferred.

hartnn (hartnn):

then i think, u need to remember atleast these 2 : log 2 = 0.301, log 3 = 0.477

OpenStudy (anonymous):

hey hw to find log of 20/5?

hero (hero):

I'm still trying to figure out the notation. What country are you from?

hero (hero):

You can interpret that in different ways: \[\log_{20}{(5)}\] \[\log_{10}(20/5)\] \[\log_{5}(20)\]

hero (hero):

If it's \[\log_{10}(20/5) = \log_{10}(4)\] , then it is relatively easy.

hero (hero):

But if it's something other than that, then idk

hero (hero):

@ruchi.

hero (hero):

If you want help, you will need to clarify what the notation is.

hero (hero):

Okay, so if it is what I think it is, then log(20/5) = 2 log(2) = .602

OpenStudy (anonymous):

hey hw u fing it?

hartnn (hartnn):

\(\log (20/5)=log 4 = log (2^2)=2log2 \\ \text{from the property of log that :} \\ \huge logx^n=nlogx\) and as i mentioned, u should remember log 2 =0.301